6 principal maxima are observed in the diffraction pattern due to a diffraction grating. Calculate the number of lines per cm in the grating if the wavelength of the light is 5000 Å.


1
Expert's answer
2021-06-07T09:29:29-0400

"(a+b)sinθ=nλ \\
N=\frac{1}{a+b} \\
θ=90° \\
(a+b)= nλ \\
(a+b) = 6 \times 5 \times 10^{-10} \\
= 3 \times 10^{-9} \\
N= \frac{1}{3 \times 10^{-9}}=3.33 \times 10^8 \;lines \; per \;m \\
= 3.33 \times 10^{10}\;lines \;per\;cm"